OpenStudy (rane):

a lovesick lad wants to throw a bag of candy and love notes into the open window of his friends bedroom 10m above he throws the gift at 60 degrees to the ground: at wht velocity should he throw the bag?

OpenStudy (rane):

@ParthKohli

Parth (parthkohli):

Let's assume that the velocity is \(x\). What is the vertical velocity of the bag of candy?

Parth (parthkohli):

You know what "vertical velocity" means, yes?

OpenStudy (rane):

yes i do know wht it means :P we r not given any velocity

Parth (parthkohli):

Yes, assume the velocity as a variable.

Parth (parthkohli):

If the velocity is \(x\), then the vertical velocity is \(v\sin(60^{\circ})\)

Parth (parthkohli):

Now, do you know what the max height is, in a projectile?

OpenStudy (rane):

in general ?\or about that q;?

Parth (parthkohli):

In general... we'll apply that to the question.

OpenStudy (rane):

no

Parth (parthkohli):

Let's derive that for a vertical motion.

Parth (parthkohli):

\[v^2 = u^2 - 2gH\]Now, in vertical motion, the max height is when \(v = 0\).\[0^2 = u^2 - 2gH \Rightarrow H_{max} = \dfrac{u^2}{2g}\]

Parth (parthkohli):

So, the max height is\[\dfrac{u^2}{2g}\]Now, what is the initial vertical velocity?

Parth (parthkohli):

And what is the max height this projectile covers?

OpenStudy (rane):

2*9.8

Parth (parthkohli):

?

Parth (parthkohli):

\[H_{\large max}=\dfrac{u^2}{2g}\]

Parth (parthkohli):

\[H_{\rm max}= 10~ m\]\[u= x\sin(60^{\circ})\]Plug that into the formula.

OpenStudy (rane):

i got 12

Parth (parthkohli):

\[200 = \dfrac{3}{4}x^2\]

OpenStudy (rane):

wht??

Parth (parthkohli):

Hmm, let's solve it.\[10 = \dfrac{\left(x\dfrac{\sqrt{3}}{2}\right)^2}{2\times 9.8}\]

Parth (parthkohli):

Get it?

OpenStudy (rane):

nop, u just lost me with the above fraction

OpenStudy (rane):

3/2 ??????? where is that from?

Parth (parthkohli):

\[u = x\sin(60) = x \dfrac{\sqrt3}{2}\]

OpenStudy (rane):

how did u get 3/2?

Parth (parthkohli):

\[\sin(60 ) = \frac{\sqrt{3}}{2}\]

OpenStudy (rane):

oh ook....

OpenStudy (rane):

i got 258?

Parth (parthkohli):

Find the root of that.